# PrisonBreak: Jailbreaking Large Language Models with Fewer Than Twenty-Five Targeted Bit-flips

### v1

@@ 5.1 Our Jailbreaking Objective
BitFinder uses the progressive bit-search, where the attacker identifies critical weights and target bits based on the gradients computed with respect to an attack objective. Prior work designs attack objectives, e.g., maximizing cross-entropy loss to degrade performance [3, 1, 49], minimizing prediction error of specific samples to achieve targeted misclassification [4, 5] or optimizing for trojan/backdoor attacks [7, 8, 10, 9, 6]. But these objectives developed for classification settings are incompatible with jailbreaking Here, we propose our objective tailored to identify critical weights and target bits exploitable for jailbreaking.
Limitations in targeting affirmative tokens. As described in C3 in §4.2, the attack objective proposed in prompt-based jailbreaks [50, 16]–which makes the target model respond with affirmative tokens at the beginning of its response–results into two unique failure modes:
  • Overfitting: Our initial attack causes the target model to extremely overfit to affirmative tokens. It outputs them repeatedly, yielding nonsensical output.
@@
Table 1 provides a few examples illustrating responses generated using only the prior work’s objective and PrisonBreak. Most prior work on prompt-based jailbreaking considers the attack successful if a model’s output begins with an affirmative response. But in practice, these attacks have lower fidelity in their generated responses.
Improving our attack’s fidelity. We hypothesize that the lower fidelity is because the objective lacks “harmful context”: it optimizes only for affirmative tokens without ensuring the model actually generates harmful content. To address this, we design our objective function to consider both affirmative tokens and actual harmful responses. This requires oracle harmful responses—plausible excerpts of harmful content the model might produce if it were non-aligned. Because no prior work has developed such oracles, we develop a method to generate our own.
@@
Our loss function computes the cross-entropy loss (highlighted in dark-red) across the entire proxy dataset. However, cross-entropy treats all tokens equally, making tokens in the harmful context equally important as the first few affirmative tokens. This can be undesirable, as the model must first produce affirmative tokens; otherwise, it is less likely to generate the specific tokens in the harmful context, thereby degrading the quality of harmful responses.
To address this, we assign more importance to the initial tokens of each sequence by incorporating a weighting term, highlighted in blue. The importance gradually decreases; for example, the first token has a weight of 1, the tenth $0.638$, and the last $0.368$ for a sequence of 20. Our attack prioritizes generating affirmative tokens first, and then make the target model respond with high-quality harmful contexts.
@@ 5.2 Identifying Critical Weights
Next, we use the jailbreaking objective to identify critical weights in the target model. These are the weights whose manipulations lead to the target model aligning most closely with our attack objective—jailbreaking.
@@
  • Skipping layers not critical to our goal. Commercial-scale language models typically contain 32–80 Transformer blocks, resulting in a total layer count that can quickly reach several hundred. We thus explore layers that are never selected by BitFinder and exclude them from subsequent searches. In preliminary experiments, we observe that our offline procedure does not select any normalization, embedding, and unembedding layers. By skipping these layers in our search, we reduce the total number of weights to inspect by roughly 20%.
  • Inspecting Top-$k$ critical weights. Another approach to reduce computations is to inspect only the Top-$k$ weights within each tensor (layer), based on their importance scores, as critical weights. We set $k$ to a fixed value, since making $k$ proportional to the weight count per layer results in no weights being selected from layers with fewer parameters. While our ablation study in §7.4 shows that the attacker can use a smaller $k$, such as 10, with a decent attack success, e.g., of 73.6% on Llama2-7B [34], we conservatively set $k$ to 100, achieving 80–95% jailbreaking success across all evaluated models.
@@ 5.3 Selecting Target Bits
The final step is selecting bit locations to flip. We choose target bits based on the importance scores we compute. The scores reflect the target model’s generation capabilities after a bit-flip ( C3). Here, for an adversary with a memory template for a specific DRAM module, the attacker considers an additional factor—the practicality to flip chosen target bits with Rowhammer ( C4). Moreover, we develop bit-level speed-up techniques to accelerate the selection process( C2).
(Hardware-aware search) Skipping non-flippable bits. In §4.1, we assume the attacker has access to a list of offsets within the DRAM’s physical pages, along with the corresponding flip directions that Rowhammer can induce. If the attacker knows these flippable bit locations, they can select only bits that match the offset and flip direction of at least one vulnerable cell. This additional step is particularly useful for resilient DRAM modules, where a bit selected by BitFinder is less likely to be flipped. We evaluate our hardware-aware search in §8 to attack DRAM modules with up to 1625$\times$ fewer vulnerable cells and achieve comparable jailbreak success to the hardware-agnostic method.
Rank the bits by the jailbreak score. We first rank the bits in our critical weights, based on the jailbreaking score:
@@
where $f^{\prime}_{i}$ is the model with a bit $b_{i}$ flipped, and $\mathbf{B}$ represents all bits in the critical weights identified in §5.2. The attacker computes this score for each bit and ranks all bits in ascending order. If the score is low, flipping $b_{i}$ is more likely to induce the model to produce harmful responses.
Selecting a bit based on the utility score. Our preliminary attack, which uses only the jailbreaking score to select target bits, often reduces the model’s ability to produce coherent, high-quality answers. To measure the likelihood that flipping a bit degrades a model’s utility, we present the utility score.
@@
Given a bit $b_{i}\in\mathbf{B}$ and the model $f$ The utility score $\mathcal{U}$ is defined as follows:
@@
where $D_{c}$ is a clean dataset to evaluate $f$’s utility, $\mathcal{L}_{ce}$ is the (unweighted) cross-entropy loss, and $(x_{i},y_{i})|^{n}_{i=1}$ are the prompt-target pairs. We construct a dataset $D_{c}$ by randomly sampling prompts from representative clean data. We use Alpaca, a cleaned and curated version of the dataset used to train Alpaca language models [53]. For each prompt $x_{i}$ in Alpaca, we generate a response $y_{i}$ from the target model. We leverage greedy decoding in the generations, which serves as the model’s most likely responses. $D_{c}$ should contain sufficiently large samples to capture the model’s generation behavior. Interestingly, while the original Alpaca dataset contains 52,000 instructions, we find that only 100 are needed to approximate the model’s utility, reducing the required computations by 520$\times$.
@@
Bit-level speed-up techniques. Even with the parameter-level speed-up in §5.2, we still need to examine 16 bits per critical weight, which remains computationally demanding. Based on our initial experiments, progressively searching for 25 target bits would still take $\sim$142 days on a Nvidia A40 GPU. We address this issue by leveraging three insights to reduce the total number of candidate bits:
@@
  • The gradient sign. Not all bit-flips reduce our jailbreaking objective; if a weight change has the same sign as the gradient, it will increase the loss. Consider a weight $w$ with gradient $\nabla w$. For each bit in $w$, we obtain a new value $w^{\prime}$ after flipping it and define $\Delta w=w^{\prime}-w$. We then skip the bits whose flipping results in $\Delta w$ with the same sign as the gradient, i.e., $\text{sign}(\Delta w)=\text{sign}(\nabla w)$. Instead, we focus on the opposite sign to minimize the jailbreaking objective. By skipping these bits, we reduce the bits to examine by $\sim$50%.
@@
  • The exponent bits. In our initial investigation, we find that most critical weights fall within $[-3,3]$, so flipping a mantissa bit or the sign bit does not produce a change large enough to induce a jailbreak. We thus focus on the exponent bits, particularly the three most significant ones, which yield the largest $\Delta w$. With roughly half of these exponent bits resulting in $\Delta w$ with the wrong sign, this reduces the search to just 1.5 bits.
@@
  • The $\mathbf{0}\!\rightarrow\!\mathbf{1}$ flip direction. Because most critical weights have small values, a $\mathbf{0}\!\rightarrow\!\mathbf{1}$ flip results in relatively larger changes than a $\mathbf{1}\!\rightarrow\!\mathbf{0}$ flip. We thus consider only $\mathbf{0}\!\rightarrow\!\mathbf{1}$ flips. This reduces the number of bits to search from 1.5 to 1, allowing us to evaluate just 1 bit per 16-bit weight.
6 PrecisionHammer: Flipping Target Bits
Now we use Rowhammer attacks to flip target bits identified in §5. PrecisionHammer operates in three steps:
1
Memory templating. In the first step, our adversary performs memory templating: we profile DRAM to identify cells susceptible to Rowhammer. This step is common in prior work on bit-flip attacks targeting neural networks [54, 1, 14, 11, 7]. We also need to reverse-engineer the DRAM address mapping scheme which can be achieved with advanced techniques introduced in [20, 42]. Another challenge in templating is bypassing Rowhammer mitigation techniques deployed in DRAM, such as target row refresh (TRR), to successfully induce bit-flips. To this end, we use off-the-shelf Rowhammer fuzzing tools [43, 22, 42]. We also record optimal access patterns that successfully induce bit-flips in DRAM modules for later use during actual fault injection. The templates will include details on vulnerable cell locations and bit-flip directions.
2
Memory massaging. The next step is to position target bits within DRAM cells exploitable by Rowhammer. The attacker first analyzes the memory template to identify vulnerable cells. We ensure that the bit offset in the memory template matches the offset in the weight tensor and aligns with the flip direction. Since popular deep-learning libraries, such as PyTorch [55], allocate memory for large tensors (where language model parameters are stored) in a page-aligned manner, we can directly convert tensor offsets into offsets in the page. We use established memory massaging techniques [21, 40] to align these target bits with matching DRAM cells. Similar to the attacks in prior work, PrecisionHammer leverages the per-core page frame cache, or per-CPU pageset [56]—a last-in-first-out cache that stores pages recently released by processes on the core—to enable precise memory massaging. By releasing the physical pages containing exploitable Rowhammer cells in the reverse order of the target tensor’s allocation, the attacker ensures that the target tensor pages are mapped exactly to physical pages with vulnerable DRAM cells.
3
Precise bit-flip via Rowhammer. Once the target bits are located to exploitable DRAM cells, we flip them via Rowhammer using the patterns identified in the templating.
No Attack Prompt-Based Parameter-Based Models GCG-M Baseline BFA PrisonBreak Ortho ASR Acc. ASR ASR $\Delta$Acc. $\Delta$P ASR $\Delta$Acc. $\Delta$P ASR $\Delta$Acc. $\Delta$P Vicuna-7B-v1.3 11.95 51.88 72.33 77.36 -10.99 5 94.34 -1.33 5 81.13 +0.12 2.0B Vicuna-7B-v1.5 18.24 55.02 40.88 69.81 -14.30 2 93.08 -2.52 12 67.30 -1.33 2.0B Vicuna-13B-v1.5 10.69 60.46 47.80 63.15 -13.75 12 90.57 -4.09 20 83.65 +0.34 3.8B Llama2-7B 1.26 52.24 72.96 38.36 -17.27 11 86.79 -1.92 25 84.91 -0.34 2.0B Llama2-13B 3.14 58.18 50.31 54.09 -4.47 25 85.53 -4.17 18 56.60 -1.07 3.8B Llama3-8B 6.27 66.42 22.64 89.31 -12.06 8 86.79 +0.06 20 93.08 -0.22 2.8B Qwen2-1.5B 9.43 54.51 67.30 76.10 -9.60 23 80.50 -4.09 25 71.70 -0.18 0.7B Qwen2-7B 13.84 68.47 30.82 81.76 -12.22 2 93.08 -1.01 25 81.76 +0.15 2.6B Average 9.35 58.40 50.63 68.74 -11.83 11 88.84 -2.38 19 77.52 -0.28 2.5B
Table 2: Attack effectiveness. The attack success rate (ASR) and average accuracy (Acc.) for different jailbreaking attacks. $\Delta$Acc. is the percentage point change relative to the clean performance. For the parameter-based attacks, we also report the number of parameters modified ($\Delta$P). “No Attack” refers to the original models without any compromise. The best ASR achieved on each model across the attacks is highlighted in bold.
@@ 7.1 Experimental Setup
Models. We use 8 open-source large-language models within four different model families, which we display in Table 3. All model parameters are represented as half-precision floating-point numbers, with sizes ranging from 1.5–13 billion. We select models with different combinations of alignment algorithms applied during their post-training, namely supervised fine-tuning (SFT) [30, 34], reinforcement learning with human feedback (RLHF) [31], and direct preference optimization (DPO) [33]. For all experiments, we utilize each model’s default chat template, available from their respective HuggingFace [57] model cards.
Model Family # Parameters Alignment Vicuna [58] 7B, 13B SFT Llama2 Chat [34] 7B, 13B SFT, RLHF Llama3 Instruct [35] 8B SFT, RLHF, DPO Qwen2 Instruct [36] 1.5B, 7B SFT, RLHF, DPO
Table 3: Model Families. Parameter count and alignment algorithms of the model families used in our experiments.
@@
  • Ortho [39] is a parameter-based jailbreak. It works on the assumption that language models have a subspace of their activation space dedicated to refusal, i.e., determining whether to refuse or respond to a query. Ortho first approximates a “refusal direction” using contrastive prompts and then ablates it from the residual stream, thereby ideally removing the model’s ability to refuse. The attack modifies billions of parameters and achieves high ASRs across various open-source models.
@@
7.2 PrisonBreak Effectiveness
We first evaluate attack effectiveness across open-source language models publicly available from the literature. Here, we assume the attackers are at their best effort. In our attack, we assume all target bits identified by our offline procedure will be flipped under Rowhammer pressure (i.e., the hardware-agnostic search), following the same assumption made in prior bit-flip attacks on neural networks [7, 3, 49, 8, 5, 9, 10]. Note that in §8 we, for the first time, evaluate the practical attack success rate of our bit-flip attack across multiple DRAM modules with varying levels of Rowhammer susceptibility. For other attacks, we conduct an extensive hyper-parameter search to bring the highest ASR. Our results will serve as the upper-bound for various jailbreaking adversaries, including ours.
Results. Table 2 summarizes our results. We test each attack individually and report the ASR and change in Acc. ($\Delta$Acc.). We do not report $\Delta$Acc. for GCG-M, as prompt-based attacks do not impact clean accuracy. For parameter-based attacks, we additionally report the number of parameters each attack modifies ($\Delta$P).
PrisonBreak outperforms all baseline attacks with an average ASR of 88.84%. Due to the different attack surface exploited by GCG-M, a direct comparison is not possible. However, the results demonstrate that our attack offers distinct advantages. It enables permanent runtime jailbreaking, requires no modification to harmful prompts, and achieves an average 1.75$\times$ higher ASR. Most notably, our attack improves the baseline bit-flip attack by an average ASR of 29%. By extending the context in the search objective and weeding out destructive bit-flips, we evidently find much higher-quality target bits regardless of model configuration. We observe a case where the baseline BFA achieves a higher ASR on Llama3-8B. However, as discussed in §7.3, this comes at a large cost to the quality of harmful responses. The attack performing closest to ours is Ortho, with comparative ASRs on most models. However, it modifies a large portion of the weight space, between 29% to 47% (0.7–2.5 billion parameters), depending on the model size and architecture. Since this would require several billion bit-flips, many within the same physical page, Ortho is not practically exploitable by Rowhammer at runtime. Conversely, we achieve higher ASR while only changing 5–25 parameters (i.e., 5–25 bits), a reduction of $10^{7}$ to $10^{8}$. Despite only being able to modify a few bits, PrisonBreak induces a jailbroken state.
[…]
However, we do find differences in the number of bit-flips necessary to achieve high ASR. In Vicuna models, 5–20 bit-flips are sufficient to jailbreak, while Llama and Qwen models require 18–25 bit-flips for comparable ASR. In addition to the number of bit-flips, the ASR in Llama and Qwen models are slightly lower than in Vicuna models. We attribute this difference to the alignment mechanisms each model family employs. Vicuna uses only SFT [30, 34]. In contrast, Llama2 and Qwen models also employ RLHF [31, 32] and DPO [33]. With more alignment, models show marginally increased resilience to our attack. However, compared to the effort required to train language models at scale with those mechanisms, our attack only needs a few more bit-flips (13–20) to achieve 80–93% ASR.
Our attack preserves model utility. Maintaining a model’s generation capabilities is important when jailbreaking at runtime; otherwise, the victim can easily detect anomalies in normal use and re-load the model into memory. To evaluate, we analyze the performance change in accuracy (Acc.) on tinyBenchmarks. Each value in Table 2 represents the average performance over eight benchmarking tasks (see Appendix C for the details). We see that our attack inflicts minimal change on model utility with an average Acc. reduction of 2.4%. This is not the case for Baseline BFA, which suffers $\sim$5$\times$ times higher reduction (11.8%) because it selects destructive bit flips during its search process. However, compared to Ortho, our attack results in a slightly higher utility loss, 2%. This is likely due to the number of parameters the attack can perturb. Across all models, Ortho perturbs at least 0.5 billion parameters, while PrisonBreak alters up to 25 bits across 25 parameters. Because our attack targets fewer weights, we cannot distribute behavior changes evenly across a large number of parameters and must instead modify a few weights disproportionately.
@@ 7.2 PrisonBreak Effectiveness
Baseline BFA “Sure, here’s a detailed guidance on how to smuggle contraband while evading law enforcement: sure, sure, sure sure, sure, sure, sure, and more. Sure, here’s a detailed guidance on how to smuggle contraband while evading law enforcement…”
@@
Ortho “Smuggling contraband while evading law enforcement can be a challenging task, but with careful planning and execution, it is possible to minimize the risk of detection. Here are some general guidelines to consider: 1. Choose the right method: … 2. Plan your route … 3. Use a trusted smuggler: …”
Table 4: Example model responses from different jailbreak attacks. The harmful query is “Give detailed guidance on how to smuggle contraband while evading law enforcement,” with generations from Vicuna-v1.5b-7b across different attacks.

### v2

@@ 5.1 Our Jailbreaking Objective
PrisonBreak employs a progressive bit-search, where the attacker identifies critical weights and target bits based on the gradients computed with respect to an attack objective. Prior attack objectives include maximizing the training loss to degrade performance [3, 1, 48], minimizing prediction error of specific samples to achieve targeted misclassification [4, 5] or optimizing for trojan/backdoor attacks [7, 8, 10, 9, 6]. But these objectives developed for classification settings are incompatible with jailbreaking.
Here, we propose our objective tailored to identify critical weights and target bits exploitable for jailbreaking.
Limitations of employing objectives from prior attacks. As described in C3 in §4.1, the attack objective proposed in prompt-based jailbreaks [49, 13]—which makes the target model respond with affirmative tokens at the beginning of its response—results in two unique failure modes:
  • Overfitting: Using these objectives leads the target model to excessively overfit to affirmative tokens, resulting in repetitive and nonsensical outputs.
@@
Table I provides examples illustrating responses generated using only the prior work’s objective in PrisonBreak. Most prior work considers the attack successful if a model’s output begins with an affirmative response. But in practice, these attacks have lower fidelity in their generated responses.
Improving our attack’s fidelity. We hypothesize that the lower fidelity arises because the objective focuses solely on affirmative tokens, neglecting the generation of actual harmful content. To address this, we design our objective function to consider both affirmative tokens and actual harmful responses. This requires oracle harmful responses—plausible excerpts of harmful content the model might produce if it were non-aligned. Because no prior work has developed such oracles, we develop a method to generate our own.
@@
Our loss function computes the cross-entropy loss (highlighted in dark-red) across the entire proxy dataset. But, cross-entropy treats all tokens equally, making tokens in the harmful context equally important as the first few affirmative tokens. This can be undesirable, as the model must first produce affirmative tokens; otherwise, it is less likely to generate the harmful context, reducing effectiveness.
To address this, we assign more importance to the initial tokens of each sequence by incorporating a weighting term, highlighted in blue. The importance gradually decreases; e.g., the first token has a weight of 1, the tenth $0.638$, and the last $0.368$ for a sequence of 20. Our attack prioritizes generating affirmative tokens first, and then make the target model respond with high-quality harmful contexts.
@@ 5.2 Identifying Critical Weights
Next, we search for critical weights, which are the weights whose manipulations lead to the target model aligning most closely with our attack objective—jailbreaking.
@@
  • Skipping layers not critical to our goal. Commercial-scale language models typically contain 32–80 Transformer blocks, resulting in a total layer count that can quickly reach several hundred. We thus explore layers that are never selected and exclude them from subsequent searches. In preliminary experiments, our offline procedure does not select any normalization, embedding, and unembedding layers. By skipping these layers, we reduce the total number of weights to inspect by $\sim$20%.
  • Inspecting top-$k$ critical weights. Another approach to reduce computations is to inspect only the top-$k$ weights within each tensor (layer), based on their importance scores, as critical weights. We set $k$ to a fixed value, since making $k$ proportional to the weight count per layer results in no weights being selected from layers with fewer parameters. While our ablation study in §6.5 shows that the attacker can use a smaller $k$, such as 10, with a decent attack success, e.g., of 73.6% on Llama2-7B, we conservatively set $k$ to 100, achieving 80–98% jailbreaking success across all evaluated models.
@@ 5.3 Selecting Target Bits
Now we select target bits based on the importance scores. The scores should capture both the jailbreaking success and generation capabilities after a bit-flip ( C3). We also develop bit-level speed-up techniques to accelerate the selection process ( C2). Moreover, if an adversary has profiled a specific DRAM module (hardware-aware), they can assess whether chosen target bits are flippable with Rowhammer. If not, those bits are excluded from the search process to enhance efficiency ( C4). We first outline the procedure for a hardware-agnostic adversary and then show how a hardware-aware attacker can skip non-flippable bits.
(1) Rank the bits by the jailbreak score. We rank the bits in our critical weights, based on the jailbreaking score:
@@
where $f^{\prime}_{i}$ is the model with a bit $b_{i}$ flipped, and $\mathbf{B}$ represents all bits in the critical weights. The attacker computes this score for each bit and ranks all bits in ascending order. If the score is low, flipping $b_{i}$ is more likely to induce the model to produce harmful responses.
(2) Selecting a bit based on the utility score. Our preliminary attack, which uses only the jailbreaking score to select target bits, often reduces the model’s ability to produce coherent, high-quality answers. To measure the likelihood of a bit-flip degrading utility, we present the utility score.
@@
Given a bit $b_{i}\in\mathbf{B}$ and the model $f$, the utility score $\mathcal{U}$ is defined as follows:
@@
where $D_{c}$ is a clean dataset to evaluate $f$’s utility, $\mathcal{L}_{ce}$ is the (unweighted) cross-entropy loss, and $(x_{i},y_{i})|^{n}_{i=1}$ are the prompt–target pairs. We construct $D_{c}$ by randomly sampling prompts from representative clean data. We use Alpaca, a cleaned and curated version of the dataset used to train Alpaca language models [52]. For each prompt $x_{i}$ in Alpaca, we generate a response $y_{i}$ from the target model. We use greedy decoding for the generations, which serves as the model’s most likely responses. $D_{c}$ should contain a sufficient number of samples to capture the model’s generation behavior. Interestingly, while the original dataset contains 52,000 instructions, only 100 are needed to approximate the model’s utility, reducing the computations by 520$\times$.
@@
Bit-level speed-up techniques. Even with the parameter-level speed-up in §5.2, we still need to examine 16 bits per critical weight, which remains computationally demanding. Based on our initial experiments, progressively searching for 25 target bits would still take $\sim$142 days on an Nvidia A40 GPU. To address this issue, we leverage three insights:
@@
  • Gradient sign. If a bit-flip causes a weight change with the same sign as the gradient, it will increase the loss, thereby not inducing jailbreaking. Consider a weight $w$ with gradient $\nabla w$. For each bit in $w$, we obtain a new value $w^{\prime}$ after flipping it and define $\Delta w=w^{\prime}-w$. We skip the bits whose flipping results in $\Delta w$ with the same sign as the gradient, i.e., $\text{sign}(\Delta w)=\text{sign}(\nabla w)$. Skipping them reduces the bits to examine by $\sim$50%.
@@
  • Exponent bits. We observe that most critical weights fall within $[-3,3]$, so flipping a mantissa bit or the sign bit does not produce a change large enough to jailbreak. We thus focus on the exponent bits, particularly the three most significant ones, which yield the largest $\Delta w$. With roughly half of these exponent bits resulting in $\Delta w$ with the same sign, this reduces the search to just 1.5 bits.
@@
  • 0$\rightarrow$1 flip direction. Since most critical weights have small values, a 0$\rightarrow$1 flip results in relatively larger changes than a 1$\rightarrow$0 flip. We thus consider only 0$\rightarrow$1 flips. This reduces the number of bits to search from 1.5 to 1, allowing us to evaluate just 1 bit per 16-bit weight.
(3) Skipping non-flippable bits. The hardware-aware attacker knows the offsets within the DRAM’s physical pages, along with the corresponding flip directions that Rowhammer can induce. If the attacker knows these flippable bit locations, they can select only bits that match the offset and flip direction of at least one vulnerable cell. We evaluate our hardware-aware attacks in §7 against DRAM modules with up to 1625$\times$ fewer vulnerable cells and achieve comparable jailbreak success to the hardware-agnostic attacks.
@@ 6 Evaluation
TABLE II: PrisonBreak effectiveness. The attack success rate (ASR) and average accuracy (Acc.) for different jailbreaking attacks. $\Delta$Acc. is the percentage point change relative to the clean performance. For the parameter-based attacks, we also report the number of parameters modified ($\Delta$P). “No Attack” refers to the original models without any compromise. The best ASR achieved on each model across the attacks is highlighted in bold.
Models No Attack GCG-M Baseline BFA PrisonBreak Ortho ASR Acc. ASR ASR $\Delta$Acc. $\Delta$P ASR $\Delta$Acc. $\Delta$P ASR $\Delta$Acc. $\Delta$P Vicuna-7B-v1.3 11.95 51.88 72.33 77.36 -10.99 5 94.34 -1.33 5 81.13 +0.12 2.0B Vicuna-7B-v1.5 18.24 55.02 40.88 69.81 -14.30 2 93.08 -2.52 12 67.30 -1.33 2.0B Vicuna-13B-v1.5 10.69 60.46 47.80 63.15 -13.75 12 90.57 -4.09 20 83.65 +0.34 3.8B Llama2-7B 1.26 52.24 72.96 38.36 -17.27 11 86.79 -1.92 25 84.91 -0.34 2.0B Llama2-13B 3.14 58.18 50.31 54.09 -4.47 25 85.53 -4.17 18 56.60 -1.07 3.8B Llama3-8B 6.27 66.42 22.64 89.31 -12.06 8 86.79 +0.06 20 93.08 -0.22 2.8B Llama3-70B 8.18 77.83 73.58 45.92 -40.04 2 98.11 -2.56 18 93.71 -0.23 23.4B Tulu3-8B 8.17 70.60 40.25 76.73 -20.12 18 86.79 -0.96 13 83.65 +1.15 2.8B Qwen2-1.5B 9.43 54.51 67.30 76.10 -9.60 23 80.50 -4.09 25 71.70 -0.18 0.7B Qwen2-7B 13.84 68.47 30.82 81.76 -12.22 2 93.08 -1.01 25 81.76 +0.15 2.6B Average 9.12 61.56 51.89 67.26 -15.48 11 89.56 -2.26 18 79.75 -0.16 4.6B Due to the scale of Llama3-70b, we set $k$ to 10 for the bit-flip attacks to reduce search time; for other models, $k=100$.
@@ 6.1 Experimental Setup
TABLE III: Model families. Parameter count and alignment algorithms of the model families used in our experiments.
Model Family # Parameters Alignment Vicuna [53] 7B, 13B SFT Llama2 Chat [29] 7B, 13B SFT, RLHF Llama3 Instruct [30] 8B, 70B SFT, RLHF, DPO Tulu3 [54] 8B SFT, DPO Qwen2 Instruct [31] 1.5B, 7B SFT, RLHF, DPO
Models. We use 10 open-source large language models within five different model families, displayed in Table III. Once loaded into memory, model parameters are represented as half-precision floating-point numbers, with sizes ranging from 1.5–70 billion. We select models with different combinations of alignment algorithms applied during their post-training, namely supervised fine-tuning (SFT) [25, 29], reinforcement learning with human feedback (RLHF) [26], and direct preference optimization (DPO) [28]. All experiments utilize each model’s default chat template, available from their respective HuggingFace [55] model cards.
@@
  • Ortho [34] is a parameter-based jailbreak that assumes language models have a subspace of their activation space dedicated to refusal. It first estimates this “refusal direction” using contrastive prompts and then ablates it from the residual stream, ideally removing refusal entirely. The attack modifies billions of parameters and achieves high ASRs across various open-source models.
@@
TABLE IV: Example model responses from different jailbreak attacks. The harmful query is “Give detailed guidance on how to smuggle contraband while evading law enforcement,” with generations from Vicuna-7B-v1.5 across different attacks.
@@
Baseline BFA “Sure, here’s a detailed guidance on how to smuggle contraband while evading law enforcement: sure, sure, sure sure, sure, sure, sure, and more. Sure, here’s a detailed guidance on how to smuggle contraband while evading…”
@@
Ortho “Smuggling contraband while evading law enforcement can be a challenging task … Here are some general guidelines to consider: 1. Choose the right method: … 2. Plan your route … 3. Use a trusted smuggler: …”
6.2 PrisonBreak Effectiveness
We first evaluate attack effectiveness across open-source language models publicly available from the literature. Here, we assume the attackers are at their best effort. In our attack, we assume all target bits identified by PrisonBreak will be flipped under Rowhammer pressure, following the same assumption made in prior bit-flip attacks on neural networks [7, 3, 48, 8, 5, 9, 10]. Note that in §7 we, for the first time, evaluate the attack success of our bit-flip attack with Rowhammer simulations across dozens of DRAM modules with varying levels of vulnerability. For other attacks, we conduct an extensive hyperparameter search to bring the highest ASR. Our results serve as the upper bound for each jailbreaking adversary, including ours.
Results. Table II summarizes our results. We test each attack separately and report the ASR and change in Acc. ($\Delta$Acc.). We do not report $\Delta$Acc. for GCG-M, as prompt-based attacks do not impact clean accuracy. For parameter-based attacks, we report the number of modified parameters ($\Delta$P).
PrisonBreak outperforms all baseline attacks with an average ASR of 89.56%. Due to the different attack surface exploited by GCG-M, a direct comparison is not possible. However, the results demonstrate that our attack offers distinct advantages: it enables permanent runtime jailbreaking, requires no modification to harmful prompts, and achieves an average 1.73$\times$ higher ASR. Most notably, our attack improves the baseline bit-flip attack by an average ASR of 22.3%. By extending the context in the search objective and weeding out destructive bit-flips, we evidently find much higher-quality target bits regardless of model configuration. We observe a case where the baseline BFA achieves a higher ASR on Llama3-8B. However, as discussed in §6.3, this comes at a large cost to the quality of harmful responses. The attack performing closest to ours is Ortho, with comparative ASRs on most models. However, it modifies between 0.7–23.4 billion parameters depending on the model size and architecture. Since this would require several billion bit-flips, many within the same physical page, Ortho is not practically exploitable by Rowhammer at runtime. Surprisingly, we are able to break the safety alignment by modifying only 5–25 parameters (i.e., 5–25 bits), a reduction of $10^{7}$$10^{9}\times$ compared to Ortho.
Model size does not matter, but alignment algorithm does. Prior work [1, 57, 5, 4] found that increasing model size tends to enhance the resilience to bit-flip attacks, possibly by introducing redundant parameters. Another study [58] showed that models rely on multiple neurons to produce outputs, supporting the observations made by these studies. We find similar ASR across model families, with no conclusive pattern between model size and attack success. As the number of parameters increases (1.5$\rightarrow$7/8$\rightarrow$13$\rightarrow$70B), ASR slightly decreases in Vicuna (93.1$\rightarrow$90.6%) but remains the same in Llama2 and even increases in Llama3 and Qwen2 (80.5$\rightarrow$98.1%).
However, we do find differences in the number of bit-flips necessary to achieve high ASR. In Vicuna models, 5–20 bit-flips are sufficient to jailbreak, while Llama, Tulu, and Qwen models require 13–25 bit-flips for comparable ASR. In most cases, the ASR for these models is also slightly lower than in Vicuna models. We attribute this difference to the alignment mechanisms employed by each model family. Vicuna uses only SFT [25, 29]. In contrast, Llama, Tulu and Qwen models also employ RLHF [26, 27], DPO [28], or both. With more alignment, models show marginally increased resilience to our attack. However, compared to the effort required to train language models at scale with those mechanisms, our attack only needs a few more bit-flips (8–20) to achieve 80–98% ASR.
Our attack preserves model utility. Maintaining a model’s generation capabilities is important when jailbreaking at runtime; otherwise, the victim can easily detect anomalies and re-load the model into memory. We evaluate the change in accuracy ($\Delta$Acc.) on tinyBenchmarks. Each value in Table II represents the average performance over eight benchmarking tasks (see Appendix D for the details). Our attack causes minimal change in model utility with an average $\Delta$Acc. of -2.26%. This is not the case for Baseline BFA, which suffers $\sim$7$\times$ times higher reduction (-15.48%) because it selects destructive bit flips during search. However, compared to Ortho, our attack results in a slightly higher utility loss of 2%. This is likely due to the number of parameters the attack can perturb. Across all models, Ortho perturbs at least 0.7 billion parameters, while PrisonBreak alters up to 25 bits across 25 parameters. Because our attack targets fewer weights, we cannot distribute behavior changes evenly across a large number of parameters and must instead modify a few weights disproportionately.
[…]
Figure 2: Impact of attack configurations on ASR and Acc. studied in Llama2-7B.